OpenStudy (anonymous):

A car is traveling at a speed of 48 miles per hour. What is the car's speed in miles per minute? How many miles will the car travel in 20 minutes? No rounding I'm having trouble changing the the hours to minutes here. The 48 miles per hour to how many miles will the car travel in 20. I'm just a little confused here. I hate word problems TT^TT

OpenStudy (anonymous):

Would the speed be 0.8 and then would the miles traveled 16?

OpenStudy (phi):

there are 60 minutes in 1 hour if you go 48 miles in 1 hour, that means you go 48 miles in 60 minutes written as a ratio \[ \frac{ 48 \text{ miles}}{60 \text{ mins} }\] you can simplify that fraction... either in steps. e.g keep dividing top and bottom by a number that goes into each... or notice that 12 works. you get 4/5 mile/min or 0.8 miles per min if you go 0.8 miles in 1 minute you will go 16*.8 miles in 16 minutes

OpenStudy (phi):

and 20 * 0.8 = 16 miles in 20 minutes

OpenStudy (phi):

if you have a ratio miles/hour and want miles/minute you can think like this: \[ \frac{ miles}{hour} \cdot \frac{hour}{min} = \frac{ miles}{\cancel{hour}} \cdot \frac{\cancel{hour}}{min}= \frac{miles}{min}\] in other words, you want to use a conversion ratio of hours per minute you multiply by \[ \frac{1 \text{ hour}}{60 \text{ min}} \] for your problem: \[ \frac{48 \text{ miles}}{1 \text{ hour}} \cdot \frac{1 \text{ hour}}{60 \text{ min}}= \frac{48 \text{ miles}}{60 \text{ min}} = 0.8 \frac{\text{miles}}{\text{min}}\]

OpenStudy (triciaal):

@phi great explanation

OpenStudy (anonymous):

Thank you. This really helps clear things up for me